-4x+5y=20 in y-intercept form and graph
Answer by mananth(16949)

the equation in slope intercept form is y=mx+b


-4x+5y=20
Add 4x on both side
-4x+4x+5y=20+4x
5y=4x+20
divide both sides by 5
y = (4/5)x+4